OpenStudy (anonymous):

Water is poured into a conical paper cup at the rate of 3/2 in3/sec. If the cup is 6 inches tall and the top has a radius of 3 inches, how fast is the water level rising when the water is 2 inches deep? The water level is rising at a rate of what?

OpenStudy (anonymous):

lordamercy any idea what the equation for the volume of a cone is?

OpenStudy (anonymous):

(pi*r^2*h)/3

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }*\pi*r ^{2}*h\]

OpenStudy (anonymous):

yeah so we need some sort of expression for the volume in terms of \(h\) i guess we have to use similar triangles or something like that

OpenStudy (anonymous):

yea I got that r=h/3

OpenStudy (anonymous):

ok maybe i made a mistake, because it is late, but i get \(\frac{r}{h}=\frac{3}{6}=\frac{1}{2}\) making \(h=2r\)

OpenStudy (anonymous):



OpenStudy (anonymous):

i could be wrong, please let me know if i made a mistake if not we can finish fairly easily

OpenStudy (anonymous):

lets assume for the sake of argument that i am right then you have \[V=\frac{\pi r^2h}{3}\] putting \(r=\frac{h}{2}\) gives \[V(h)=\frac{\pi h^3}{12}\]

OpenStudy (anonymous):

taking derivatives gives \[V'=\frac{\pi h^2}{4}h'\]

OpenStudy (anonymous):

you are told \(V'=\frac{2}{3}\) and that \(h=2\) solve for \(h'\)

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

oh right \(V'=\frac{3}{2}\) right

OpenStudy (anonymous):

I think it's h=3r

OpenStudy (anonymous):

but we have to solve in terms of h, I think

OpenStudy (anonymous):

yeaaaa it's r=h/3

OpenStudy (anonymous):

not sure where the 3 comes from the radius is 3, the height is 6 the radius is r, the height is h that means i think that \(h=2r\) oh and i messed up

OpenStudy (anonymous):

no no i didn't we want \(h\) not \(r\) so \(r=\frac{h}{2}\)

OpenStudy (anonymous):

you may be right, but can you explain why \(r=\frac{h}{3}\) ?

OpenStudy (anonymous):



OpenStudy (anonymous):

oh goooof I can't lol sorry i divided 3 by 6 and thought it's 1/3, so sorry

OpenStudy (anonymous):

so the final answer is 3/2*pi???

OpenStudy (anonymous):

i didn't get to the final answer yet, but be can do it now

OpenStudy (anonymous):

\[V'=\frac{\pi h^2}{4}h'\] or \[h'=\frac{4V'}{\pi h^2}\]

OpenStudy (anonymous):

put \(h=2, V'=\frac{3}{2}\) and you are done

OpenStudy (anonymous):

no that is wrong

OpenStudy (anonymous):

now i think it is \[h'=\frac{3}{2\pi}\]

OpenStudy (anonymous):

Thank YOU!